OpenStudy (anonymous):

What are the first 5 terms to the pattern with the formula 21n?

OpenStudy (anonymous):

@Shay17

OpenStudy (anonymous):

thats it ?

OpenStudy (anonymous):

whats the pattern

OpenStudy (anonymous):

just whats the first 5 terms for the formula 21n its making a pattern

OpenStudy (anonymous):

idk ask @dan815

OpenStudy (anonymous):

I think you need to do 21*1 = ? 21*2 = ? 21*3 = ? 21*4 = ? 21*5 = ? But I'm not fully sure :/

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Blank  @mathmale

OpenStudy (mathmale):

In more formal terms, the formula is \[a _{n}=21n\]

OpenStudy (mathmale):

and if you're positive that the problem statement reads '21 n' then Shay's suggestion is correct. Let n begin with 1 and end with 5 to obtain the first 5 terms of this sequence.

OpenStudy (anonymous):

ok great so @Lindsay_552 do what i did

OpenStudy (anonymous):

ok thanks both of u

OpenStudy (anonymous):

@dan815 you late

OpenStudy (dan815):

mathmale is a professor!

OpenStudy (dan815):

He's an awesome teacher hes helped me in the past

OpenStudy (mathmale):

@dan815: :) Thank you so much.

OpenStudy (dan815):

He's an awesome teacher. He's helped me in the past! (I must clean up my grammar when complimenting a prof! hehe)